15x^4 - 13x^2 + 2 = 0
Solving equations in quadriatic form
please help if you can.

Found 2 solutions by dolly, stanbon:
Answer by dolly(163)


15x^4 - 13x^2 + 2 = 0
15x^4 - 10 x^2 - 3x^2 + 2 = 0
==> 5x^2(3x^2 - 2) -1(3x^2 - 2) = 0
==> (3x^2 - 2) (5x^2 - 1) = 0

==> 3x^2 - 2 = 0 or 5x^2 - 1 = 0
==> x^2 = 2/3 or x^2 = 1/5
x = +/- sqrt(2/3) or x = +/- sqrt(1/5)

Answer by stanbon(57342)

15x^4 - 13x^2 + 2 = 0
(15x^4 -10x^2 -3x^2 +2) = 0
5x^2(3x^2-2) -(3x^2-2) = 0
(3x^2-2)(5x^2-1) = 0
x^2 = 2/3 or x^2=1/5
x=+-(sqrt6)/3 or x=+-(sqrt5)/5
===================
Cheers,
Stan H.